find the second deravstive f(x)= ln (5x+1)
Answer #1

). . t 52+1 5041 Here f(x)= an (5ntly diff. W. se to x, p(x)a (5) = f(x) = 5 again diff. w. stor, gil(a) = 5. (- Snja (5) a